How many litres of 70% alcohol needs to be added to 50litres of 40% alcohol to make a 50% solution.
Answer by richwmiller(17219)

We have 50 liter of 40% alcohol
We want to add some 70% alcohol to make 50% alcohol
0.7*x+0.4*50=0.5(50+x)
0.7*x+20=25+0.5x
0.7x-0.5x=25-20
0.2x=5
x=25 liter at 70% (17.5) is added to 50 liter at 40%(20) to make 50% of 75 liter(37.5)
check
0.7*25+0.4*50=0.5(50+25)
17.5+20=0.5(75)
37.5=37.5
ok